mrs. carter invested a sum of money at 4 %. she invested a second sum, $250 more than the first sum,at 6%. if the total annual income was $90, how much did she invest at each rate?
Answer by checkley79(3341)

.04x+.06(x+250)=90
.04x+.06x+15=90
.10x=90-15
.10x=75
x=75/.10
x=$750 amount invested @ 4%
750+250=$1,000 invested @ 6%
Proof:
.04*750+.06*1,000=90
30+60=90
90=90